OpenStudy (anonymous):

Find two values of k so that each trinomial can be factored over the integers. 12x^2+kx+14

OpenStudy (raden):

just, hint first. D = m^2 wth D = discriminant and m integer

OpenStudy (raden):

169 is works :)

OpenStudy (raden):

12x^2+kx+14 D = m^2 k^2 - 4(12)(14) = m^2 k^2 - 672 = m^2 k^2 - m^2 = 672 case I : (k+m)(k-m) = 336 * 2 k+m = 336 k-m = 2 --------- (+) 2k = 338 k = 169 k-m = 2 169-m=2 m=167 case II : k^2 - m^2 = 672 (k+m)(k-m) = 168 * 4 k+m = 168 k-m=4 -------- (+) 2k = 172 k = 86 so, k = 86 is works too for the others, maybe there are many possibles for k...

OpenStudy (anonymous):

This way may be simpler. The 2 values of k are : k = -26 and k = 26. a) k = -26. This makes a + b + c = 0 (Tip 1). In this case, one real root is (1) and the other id (c/a = 14/12) y = (x -1) (12x - 14) b) k = 26. This make a - b + c = 0. One real root is (-1) and the other is (-c/a = -14/12) y = (x + 1)(12x + 14) REMINDER. Quick solving quadratic equations a) Tip 1. When a + b + c = 1, one real root is (1) and the other is (c/a) b) Tip 2. When a - b + c = 0, one real root is (-1) and the other is (-c/a). Remember these 2 TIPs. They will save you lot of time.

OpenStudy (raden):

good idea :)